OpenStudy (anonymous):

There are 20 students in a class. 12 are girls, 8 are boys. 3 girls like to play softball, 4 boys like to play softball. Find the probability of picking two students from the class and choosing a girl or a softball player. Explain your work.

OpenStudy (anonymous):

Girls 12/20 and Softball Players: 7/20

OpenStudy (anonymous):

Since this is a dependent event, shouldn't I multiply both 12/20 and 7/20 to get 21/100?

OpenStudy (anonymous):

A = choosing a girl B = choosing a softball player since they are not mutually exclusive events: P(AorB) = P(A) + P(B) - P(AandB) = P(A) + P(B) - P(A)P(B)

OpenStudy (phi):